OpenStudy (anonymous):

Please help! Medal will be given Y=3x-5;(0,5),(-1,-8)

OpenStudy (anonymous):

for (0,5) i think it must be (0,-5) 5=3(0)-5 for (-1,-8) -8=3(-1)-5

OpenStudy (nincompoop):

what do you want us to do with your question?

OpenStudy (mathmale):

Brittnay, it'd be clearer for you as well as for potential helpers if you'd please include instructions when you post a problem. What's the objective here?

OpenStudy (nincompoop):

do you even understand what the question is?

OpenStudy (mathmale):

See...nincompoop sees this matter in the same way as I.

OpenStudy (nincompoop):

does the line with an equation Y=3x-5 passes through the points (0,5),(-1,-8)

OpenStudy (nincompoop):

one way to approach this matter is to find your slope from the given two coordinates (0,5),(-1,-8) m = (y2-y1)/(x2-x1) assign your coordinates (0, 5) as x1 = 0 and y1 = 5 (-1, -8) as x2 = -1 and y2 = -8 now your turn solve for m